# FiRe: Fixed-points of Restoration Priors for Solving Inverse Problems

### v1

@@ Abstract
Selecting an appropriate prior to compensate for information loss due to the measurement operator is a fundamental challenge in imaging inverse problems. Implicit priors based on denoising neural networks have become central to widely-used frameworks such as Plug-and-Play (PnP) algorithms. In this work, we introduce Fixed-points of Restoration (FiRe) priors as a new framework for expanding the notion of priors in PnP to general restoration models beyond traditional denoising models. The key insight behind FiRe is that natural images emerge as fixed points of the composition of a degradation operator with the corresponding restoration model. This enables us to derive an explicit formula for our implicit prior by quantifying invariance of images under this composite operation. Adopting this fixed-point perspective, we show how various restoration networks can effectively serve as priors for solving inverse problems. The FiRe framework further enables ensemble-like combinations of multiple restoration models as well as acquisition-informed restoration networks, all within a unified optimization approach. Experimental results validate the effectiveness of FiRe across various inverse problems, establishing a new paradigm for incorporating pretrained restoration models into PnP-like algorithms.
@@ 1 Introduction
where $A$ is a linear operator, modeling the sensing system, and $e\in\mathbb{R}^{m}$ is some additive noise. A standard approach for solving such problem consists in finding an estimate $x$ that maximises the posterior distribution $p(x|y)$. Assuming that the noise $e$ is Gaussian, the Bayes’ rule yields the following equivalent minimization problem:
@@
From a practical perspective, PnP methods alternate between data-fidelity steps enforcing consistency with (1) and denoising steps that implicitly encode natural image priors. Despite theoretical convergence results ensuring stable reconstructions [33, 20, 11, 28], achieving state-of-the-art results relies heavily on heuristics such as decreasing step sizes or early stopping [31, 32, 36]. This gap between theory and practice stems from a critical limitation of PnP algorithms: at convergence, denoisers trained on Gaussian-corrupted images are applied to natural images, creating a significant distribution mismatch.
Figure 1: (a) Experimental setup. Two restoration models trained for different tasks: $\operatorname{R}_{1}$ is the SCUNet from [37] for JPEG restoration, and $\operatorname{R}_{2}$ is a Restormer model [34] fine-tuned for Gaussian deblurring. (b) Fixed points of restoration models are not natural images. Direct iteration of these models in a standard PnP framework yields strong artifacts. (c) Composition with training degradation preserves natural image characteristics. For each model $R_{i}$ and its associated degradation $D_{i}=H_{i}\cdot+w_{i}$, the fixed-points of $\operatorname{R}_{i}\circ D_{i}$ maintain natural image properties. (d) Linear combinations yield improved fixed points. Combining $\operatorname{R}_{1}\circ D_{1}$ and $\operatorname{R}_{2}\circ D_{2}$ produces higher quality fixed-points (top), with stable PSNR values (bottom graph, solid lines) compared to the rapidly degrading naive approach from (b) (bottom graph, dashed lines).
A natural extension of this framework is to consider other restoration models as potential implicit priors. Recent works demonstrate that super-resolution models can be incorporated into both PnP restoration [9] and diffusion models [2]. This approach is particularly appealing as it can leverage the vast collection of pretrained image restoration models available to the community. However, this approach is not as straightforward as substituting a denoiser with a more general restoration model, as this naive approach leads to poor reconstruction quality. Moreover, unlike for denoising models, no equivalent of Tweedie’s formula has been established for general restoration models.
To leverage the implicit prior of restoration models in the wild, our main observation is that both denoisers and restoration models are trained to have natural images as fixed points of the composition between the train-time image degradation (e.g. additive noise or linear degradation) and the restoration model. Based on this observation, we propose the Fixed-points of Restoration priors (FiRe) framework, which enables using any restoration model coupled with the degradation it was trained on as an implicit prior to solve (1).
@@ 2.1 Denoising priors and PnP algorithms
While early methods proposed explicit priors derived from signal-processing arguments (e.g. sparsity in a wavelet basis [23] or of the gradient image [3]), recent progress in imaging inverse problems relies on implicit denoising-based priors and their connection to score-matching [14, 7]. Minimum Mean Square Error (MMSE) denoisers play a crucial role to implicitly probe the prior distributions of natural images. For a given noisy sample $y=x+n$ where $x\sim p(\text{data})$ and $n\sim\mathcal{N}(0,\sigma^{2})$, these operators $\operatorname{R}$ satisfy $\operatorname{R}(y)=\mathbb{E}[x|y]$. Starting from this conditional expectation, Tweedie’s formula yields [33, 16]
@@
thus providing a link between the denoiser $\operatorname{R}$ and the sought prior. In turn, assuming that $-\nabla\log p$ is well approximated by a neural network $\operatorname{R}$, the popular RED approach suggests the following iteration
@@
While these approaches have shown promising results, the methods in (4) and (5) often yield suboptimal reconstructions. We argue that this limitation stems from two key issues: first, during inference, the denoiser processes inputs that deviate significantly from its training distribution of Gaussian noise-corrupted images. This mismatch is particularly problematic at convergence, where the denoiser’s action must align with the score of the inverse problem’s noise to yield stable solutions. Second, and more fundamentally, natural images are not fixed-points of denoising networks: even for simple denoising problems where $A=\operatorname{Id}$, iterating $\operatorname{R}$ in (5) leads to artifacts, similar to those shown in Figure 1 (b). In this work, we adopt the view-point of the fixed-point sequence generated by (4) to overcome the shortcomings of this approximate Tweedie formula as well as the distribution shift at inference time.
@@ 2.2 Related works
The link between denoising autoencoders and score matching was first established in [30], showing that denoising autoencoders implicitly estimate the score $\nabla\log p(x)$ of the data distribution $p(x)$. Implicit denoising priors were introduced by [29], who proposed replacing classical proximity operators with denoising operators in ADMM algorithms, interpreting the denoising operation as a MAP estimator. This framework was later revisited by [33], who formalized the connection through Tweedie’s formula (3) for MMSE denoisers and provided a proximal operator interpretation. Subsequent work by [16] deepened the theoretical understanding of the underlying distributions and proposed a PnP version of Langevin sampling. Alternative viewpoints emerged through the study of monotone operators [25, 20], and modifications of the denoising neural network architectures for nonconvex priors [11]. Our theoretical analysis diverges from previous approaches that rely on Lipschitz properties [20, 11, 33], instead building upon fixed-point arguments [6].
@@ 3.1 Fixed points of restoration models
A central observation to our work is that image restoration networks inherently possess a fixed-point property with respect to natural images. More precisely, consider a general restoration problem $y=Hx+w$, where $H$ is a linear operator and $w$ denotes additive noise. In the standard supervised setting, image restoration models are trained by minimizing an objective of the form
@@
In this assumption, we have used the following definition: a closed set $C$ is prox-regular if each point in $C$ has a neighborhood where the projection onto $C$ is single-valued [17]. Although the prox-regularity of $C$ might be too restrictive for natural images, this assumption is weaker than convexity and it is satisfied by a broad class of sets including smooth submanifolds of $\mathbb{R}^{n}$ [21]. The following result, from [17], allows to link the projection on the set $C$ to the gradient of a smooth function.
@@
In practice, restoration models only approximately minimize (6) over a set of natural images. Therefore, the fixed point set associated to one restoration model may contain slightly degraded images, or natural images might only be in its neighborhood. To overcome this difficulty, we propose to rely on an ensemble of restoration priors. More precisely, for a given restoration model $\operatorname{R}_{\xi}$ and degradation $D_{\xi}$, we denote $C_{\xi}$ the associated fixed-point set. While $C_{\xi}$ may contain spurious fixed-points due to the approximate nature of the training, these spurious elements are unlikely to be shared with sets $C_{\xi^{\prime}}$ associated with models $\operatorname{R}_{\xi^{\prime}}$ trained for different tasks $D_{\xi^{\prime}}$. On the contrary, we expect that natural images should lie in the neighbourhood of all these sets. By considering the intersection $\bigcap_{\xi}C_{\xi}$, we can obtain a better approximation of the set $C$ containing natural images, as this intersection filters out model-specific artifacts. Following Section 3.1, this naturally leads to considering the log-prior $\mathbb{E}_{\xi}[d_{C_{\xi}}^{2}(x)]$, which measures how far an image $x$ is from the intersection of the sets $C_{\xi}$. With this log-prior, even if the sets’ intersection is empty, the maximum likelihood points lie the closest possible to each of these sets.
@@ 5.1 Considered priors & baselines
Category Backbone Restoration Prior Gaussian Blur Motion Blur SR×$\times$4
none - 21.57 19.48 21.69
PnP DRUNet Denoising 25.93 27.95 22.54
SCUNet Blind Denoising 23.91 30.19 23.05
FiRe LAMA Random Inpainting 24.72 26.19 21.80
LAMA Random Brush Inpainting 24.37 24.35 21.83
Restormer Gaussian Deblurring 27.01 29.83 24.29
Restormer Motion Deblurring 25.92 29.17 21.17
SCUNet Blind Denoising 28.03 30.84 24.30
SCUNet JPEG Restoration 27.19 31.23 24.01
SwinIR SR×$\times$2 28.19 27.17 24.02
SwinIR SR×$\times$3 27.31 26.84 24.77
Table 1: PSNR values on the Set3C dataset for different restoration problems. The first row (“Baseline”) corresponds to the pseudo-inverse reconstruction baseline. The second category (“PnP”) shows results from a standard Plug-and-Play (PnP) approach using DRUNet and SCUNet. The third category (“FiRe”) presents results using different restoration models and training tasks. Best results per column are in bold, second best underlined.
@@ 5.2 Single restoration prior
$y$
PSNR
PnP FiRe
Refer to caption Refer to caption
25.45 dB 25.96 dB
@@
Figure 2: Reconstruction results with various algorithms for a SR×$\times$4 problem. Left: reconstructions obtained with the SCUNet JPEG prior with the PnP approach and proposed FiRe approach. Right: (d)-(h) reconstructions obtained with Algorithm 1 for various priors.
We first evaluate Algorithm 1 in its simplest form using a single restoration prior ($N=1$). Table 1 reports PSNR values on the Set3C dataset for various restoration models and their associated training tasks. For comparison, we include two PnP baselines: the DRUNet denoising model and the blind SCUNet model. The PnP framework is recovered within Algorithm 1 by setting $H_{k}=\operatorname{Id}$ and $w_{k}=0$. Notably, most restoration priors within our FiRe framework achieve comparable or better performance than the DRUNet baseline, with the exception of LAMA inpainting priors (which still outperform the pseudo-inverse reconstruction). Also note that the SCUNet blind denoiser performs significantly better within our FiRe framework than in the PnP setting. This suggests that our fixed-point formulation successfully generalizes the denoising prior paradigm to a wide class of restoration models. Figure 2 (right panel) shows the restoration result for the SCUNet JPEG prior in a PnP vs FiRe framework. Note how the latter reconstruction has less artifacts.
Associated visual for the SR×$\times 4$ problem are shown in Figure 2 (left panel). We notice that except the LAMA inpainting priors, all priors perform on par with DPIR, and while all priors seem to yield similar reconstructions, few stand out. The LAMA prior seems to introduce glitch-like artifacts in the reconstruction, while the motion deblurring prior seems to introduce motion-like artifacts. All other priors seem to slightly over-smooth the reconstruction. The fact that the LAMA inpainting prior enables meaningful restoration is surprising, since it was only trained on binary mask inpainting. In particular, setting all mask values to one would reduce the problem to simple least squares minimization (i.e., pseudo-inverse restoration), whereas introducing sparse masks enables the model to leverage its learned prior. The model’s inability to remove noise, which we attribute to its noiseless training regime, further demonstrates how the training setup influences the prior’s behavior. This observation supports our broader claim that restoration models can impose meaningful priors even when trained on tasks significantly different from the target problem.
5.3 Combining multiple priors
@@ 5.4 Conditioning the prior on the measurements
(a) Observed (b) Denoise First (c) Inpaint First (d) Denoising prior (e) Inpainting prior
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(f) DiffPIR (g) DRP [9] (h) ShaRP [10] (i) Proposed
Refer to caption Refer to caption Refer to caption Refer to caption
To showcase this strategy, we consider a noisy image inpainting problem, where $A$ is a binary masking operator. We implement Algorithm 1 with two complementary priors: an inpainting-specific model (LAMA) as $\operatorname{R}_{1}$ with $H_{1}=A$, and a general denoising model (SCUNet) as $\operatorname{R}_{2}$. Visual results in Figure 5 show that while neither $\operatorname{R}_{1}$ nor $\operatorname{R}_{2}$ alone succeeds in our FiRe framework (panels (d) and (e)), their combination successfully addresses both the inpainting and denoising aspects of the problem (panel (i)).

### v2

@@ Abstract
Selecting an appropriate prior to compensate for information loss due to the measurement operator is a fundamental challenge in imaging inverse problems. Implicit priors based on denoising neural networks have become central to widely-used frameworks such as Plug-and-Play (PnP) algorithms. In this work, we introduce Fixed-points of Restoration (FiRe) priors as a new framework for expanding the notion of priors in PnP to general restoration models beyond traditional denoising models. The key insight behind FiRe is that smooth images emerge as fixed points of the composition of a degradation operator with the corresponding restoration model. This enables us to derive an explicit formula for our implicit prior by quantifying invariance of images under this composite operation. Adopting this fixed-point perspective, we show how various restoration networks can effectively serve as priors for solving inverse problems. The FiRe framework further enables ensemble-like combinations of multiple restoration models as well as acquisition-informed restoration networks, all within a unified optimization approach. Experimental results validate the effectiveness of FiRe across various inverse problems, establishing a new paradigm for incorporating pretrained restoration models into PnP-like algorithms. Code available at https://github.com/matthieutrs/fire.
@@ 1 Introduction
where $A$ is a linear operator, modeling the sensing system, and $e\in\mathbb{R}^{m}$ is some additive noise. A standard approach for solving such problem consists in finding an estimate $x$ that maximizes the posterior distribution $p(x|y)$. Assuming that the noise $e$ is Gaussian, the Bayes’ rule yields the following equivalent minimization problem:
@@
From a practical perspective, PnP methods alternate between data-fidelity steps enforcing consistency with (1) and denoising steps that implicitly encode natural image priors. Despite theoretical convergence results ensuring stable reconstructions [36, 22, 12, 31], achieving state-of-the-art results relies heavily on heuristics such as decreasing step sizes or early stopping [34, 35, 39]. This gap between theory and practice stems from a critical limitation of PnP algorithms: at convergence, denoisers trained on Gaussian-corrupted images are applied to smooth images, creating a significant distribution mismatch.
Figure 1: (a) Experimental setup. Two restoration models trained for different tasks: $\operatorname{R}_{1}$ is the SCUNet from [40] for JPEG restoration, and $\operatorname{R}_{2}$ is a Restormer model [37] fine-tuned for Gaussian deblurring. (b) Fixed points of restoration models are not smooth images. Direct iteration of these models in a standard PnP framework yields strong artifacts. (c) Composition with training degradation preserves image characteristics. For each model $\operatorname{R}_{i}$ and its associated degradation $D_{i}=H_{i}\cdot+w_{i}$, the fixed-points of $\operatorname{R}_{i}\circ D_{i}$ maintain natural image properties. (d) Linear combinations yield improved fixed points. Combining $\operatorname{R}_{1}\circ D_{1}$ and $\operatorname{R}_{2}\circ D_{2}$ produces higher quality fixed-points (top), with stable PSNR values (bottom graph, solid lines) compared to the rapidly degrading naive approach from (b) (bottom graph, dashed lines).
A natural extension of this framework is to consider other restoration models as potential implicit priors. Recent works demonstrate that super-resolution models can be incorporated into both PnP restoration [10] and diffusion models [2]. This approach is particularly appealing as it can leverage the vast collection of pretrained image restoration models available to the community. However, this approach is not as straightforward as substituting a denoiser with a more general restoration model, as this naive approach leads to poor reconstruction quality. Moreover, extending Tweedie’s formula to general restoration models beyond denoisers is not trivial.
To leverage the implicit prior of restoration models in the wild, our main observation is that both denoisers and restoration models are trained to have smooth images as fixed points of the composition between the train-time image degradation (e.g. additive noise or linear degradation) and the restoration model. Based on this observation, we propose the Fixed-points of Restoration priors (FiRe) framework, which enables using any restoration model coupled with the degradation it was trained on as an implicit prior to solve (1).
@@ 2.1 Denoising priors and PnP algorithms
While early methods proposed explicit priors derived from signal-processing arguments (e.g. sparsity in a wavelet basis [26] or of the gradient image [4]), recent progress in imaging inverse problems relies on implicit denoising-based priors and their connection to score-matching [15, 8]. Minimum Mean Square Error (MMSE) denoisers play a crucial role to implicitly probe the prior distributions of smooth images. For a given noisy sample $y=x+n$ where $x\sim p(\text{data})$ and $n\sim\mathcal{N}(0,\sigma^{2})$, these operators $\operatorname{R}$ satisfy $\operatorname{R}(y)=\mathbb{E}[x|y]$. Starting from this conditional expectation, Tweedie’s formula yields [24, 36, 17]
@@
thus providing a link between the denoiser $\operatorname{R}$ and the sought prior. In turn, defining $f(x)=\frac{1}{2}\|Ax-y\|^{2}$, and assuming that $-\nabla\log p$ is well approximated by a neural network $\operatorname{R}$, the popular RED approach suggests the following iteration
@@
While these approaches have shown promising results, the methods in (4) and (5) often yield suboptimal reconstructions. We argue that this limitation stems from two key issues: first, during inference, the denoiser processes inputs that deviate significantly from its training distribution of Gaussian noise-corrupted images. This mismatch is particularly problematic at convergence, where the denoiser’s action must align with the score of the inverse problem’s noise to yield stable solutions. Second, and more fundamentally, smooth images are not fixed-points of denoising networks: even for simple denoising problems where $A=\operatorname{Id}$, iterating $\operatorname{R}$ in (5) leads to artifacts, similar to those shown in Figure 1 (b). In this work, we adopt the view-point of the fixed-point sequence generated by (4) to overcome the shortcomings of this approximate Tweedie formula as well as the distribution shift at inference time.
@@ 2.2 Related works
The link between denoising autoencoders and score matching was first established in [33], showing that denoising autoencoders implicitly estimate the score $\nabla\log p(x)$ of the data distribution $p(x)$. Implicit denoising priors were introduced by [32], who proposed replacing classical proximity operators with denoising operators in ADMM algorithms, interpreting the denoising operation as a MAP estimator. This framework was later revisited by [24, 36], who formalized the connection through Tweedie’s formula (3) for MMSE denoisers and provided a proximal operator interpretation. Subsequent work by [17] deepened the theoretical understanding of the underlying distributions and proposed a PnP version of Langevin sampling. Alternative viewpoints emerged through the study of monotone operators [28, 22], and modifications of the denoising neural network architectures for nonconvex priors [12]. Our theoretical analysis diverges from previous approaches that rely on Lipschitz properties [22, 12, 36], instead building upon fixed-point arguments [7].
@@ 3.1 Fixed points of restoration models
A central observation to our work is that image restoration networks inherently possess a fixed-point property with respect to smooth images. More precisely, consider a general restoration problem $y=Hx+w$, where $H$ is a linear operator and $w$ denotes additive noise. In the standard supervised setting, image restoration models are trained by minimizing an objective of the form
@@
In this assumption, we have used the following definition: a closed set $C$ is prox-regular if each point in $C$ has a neighborhood where the projection onto $C$ is single-valued [18]. Although the prox-regularity of $C$ might be too restrictive for smooth images, this assumption is weaker than convexity and it is satisfied by a broad class of sets including smooth submanifolds of $\mathbb{R}^{n}$ [23]. The following result, from [18], allows to link the projection on the set $C$ to the gradient of a smooth function.
@@
In practice, restoration models only approximately minimize (6) over a set of natural images. Therefore, the fixed point set associated to one restoration model may contain slightly degraded images, or smooth images might only be in its neighborhood. To overcome this difficulty, we propose to rely on an ensemble of restoration priors. More precisely, for a given restoration model $\operatorname{R}_{\xi}$ and degradation $D_{\xi}$, we denote $C_{\xi}$ the associated fixed-point set. While $C_{\xi}$ may contain spurious fixed-points due to the approximate nature of the training, these spurious elements are unlikely to be shared with sets $C_{\xi^{\prime}}$ associated with models $\operatorname{R}_{\xi^{\prime}}$ trained for different tasks $D_{\xi^{\prime}}$. On the contrary, we expect that smooth images should lie in the neighborhood of all these sets. By considering the intersection $\bigcap_{\xi}C_{\xi}$, we can obtain a better approximation of the set $C$ containing smooth images, as this intersection filters out model-specific artifacts. Following Section 3.1, this naturally leads to considering the log-prior $\mathbb{E}_{\xi}[d_{C_{\xi}}^{2}(x)]$, which measures how far an image $x$ is from the intersection of the sets $C_{\xi}$. With this log-prior, even if the sets’ intersection is empty, the maximum likelihood points lie the closest possible to each of these sets.
@@ 5.1 Considered priors & baselines
Backbone Restoration prior Gaussian Blur Motion Blur SR×$\times$4
none - 21.57 19.48 21.69
PnP DRUNet denoising 25.93 27.95 22.54
SCUNet blind denoising 23.91 30.19 23.05
ShaRP LAMA random inpainting 23.42 23.27 21.96
Restormer gaussian debluring 27.14 28.12 23.16
SwinIR SR×$\times$2 26.01 26.66 23.24
FiRe LAMA random inpainting 24.72 26.19 21.80
random brush inpainting 24.37 24.35 21.83
Restormer gaussian deblurring 27.01 29.83 24.29
motion deblurring 25.92 29.17 21.17
SCUNet blind denoising 28.03 30.84 24.30
JPEG restoration 27.19 31.23 24.01
SwinIR SR×$\times$2 28.19 27.17 24.02
SR×$\times$3 27.31 26.84 24.77
Table 2: PSNR values on the Set3C dataset for different restoration problems. The first row (“none”) corresponds to the pseudo-inverse reconstruction baseline. The second row shows results from a standard PnP approach using DRUNet. Third row shows results with the ShaRP algorithm. Subsequent rows present results from Algorithm 1 using different restoration models (LAMA, Restormer, SCUNet, SwinIR) and their associated training tasks. Best results per column in bold, second best underlined.
@@ 5.2 Single restoration prior
We first evaluate Algorithm 1 in its simplest form using a single restoration prior ($N=1$). Table 2 reports PSNR values on the Set3C dataset for various restoration models and their associated training tasks. For comparison, we include two PnP baselines: the DRUNet denoising model and the blind SCUNet model. The PnP framework is recovered within Algorithm 1 by setting $H_{k}=\operatorname{Id}$ and $w_{k}=0$. Most restoration priors within our FiRe framework achieve comparable or better performance than the DRUNet baseline, with the exception of LAMA inpainting priors (which still outperform the pseudo-inverse reconstruction). Also note that the SCUNet blind denoiser performs significantly better within our FiRe framework than in the PnP setting. This suggests that our fixed-point formulation successfully generalizes the denoising prior paradigm to a wide class of restoration models. We also compare to ShaRP using the same restoration priors. To achieve good performance with ShaRP required increasing the iteration limit to 200 (vs. 30 for FiRe), significantly raising the computational cost. Figure 2 shows the restoration result for the SCUNet JPEG prior in a PnP vs FiRe framework. Note how the latter reconstruction has less artifacts.
Associated visual for the SR×$\times 4$ problem are shown in Figure 3. We notice that except the LAMA inpainting priors, all priors perform on par with DPIR, and while all priors seem to yield similar reconstructions, few stand out. The LAMA prior seems to introduce glitch-like artifacts in the reconstruction, while the motion deblurring prior seems to introduce motion-like artifacts. All other priors seem to slightly over-smooth the reconstruction. The fact that the LAMA inpainting prior enables meaningful restoration is surprising, since it was only trained on binary mask inpainting. In particular, setting all mask values to one would reduce the problem to simple least squares minimization (i.e., pseudo-inverse restoration), whereas introducing sparse masks enables the model to leverage its learned prior. The model’s inability to remove noise, which we attribute to its noiseless training regime, further demonstrates how the training setup influences the prior’s behavior. This observation supports our broader claim that restoration models can impose meaningful priors even when trained on tasks significantly different from the target problem.
We provide in Fig. 10 results showing the influence of the degradation strength on the reconstruction quality. We observe that the strength of the degradation plays a role similar to that of the regularization parameter in traditional variational methods.
$y$ PnP FiRe
Refer to caption Refer to caption Refer to caption
PSNR 25.45 dB 25.96 dB
Figure 2: Reconstruction results on the SR×$\times$4 problem with the SCUNet JPEG prior.
@@
Figure 3: Reconstruction results with various algorithms for a SR×$\times$4 problem. (d)-(h) show reconstructions obtained with Algorithm 1 for various priors.
@@ 5.4 Conditioning the prior on the measurements
(a) Observed (b) Denoising prior (c) Inpainting prior
Refer to caption Refer to caption Refer to caption
(d) DRP [10] (e) Denoise first (f) Inpaint First
Refer to caption Refer to caption Refer to caption
(g) ShaRP [11] (h) DiffPIR (i) Proposed
Refer to caption Refer to caption Refer to caption
@@
To showcase this strategy, we consider a noisy image inpainting problem, where $A$ is a binary masking operator. We implement Algorithm 1 with two complementary priors: an inpainting-specific model (LAMA) as $\operatorname{R}_{1}$ with $H_{1}=A$, and a general denoising model (SCUNet) as $\operatorname{R}_{2}$. Visual results in Figure 6 show that while neither $\operatorname{R}_{1}$ nor $\operatorname{R}_{2}$ alone succeeds in our FiRe framework (panels (b) and (c)), their combination successfully addresses both the inpainting and denoising aspects of the problem (panel (i)).